(1)
A car travels around a bend of radius 400m on a road which is banked at an angle to the horizontal, if the car has no tendency to slip when travelling at 35m/s. Find the angle ..
1
Expert's answer
2019-11-11T15:51:00-0500

According to Newton’s second law of motion:

"F=m*a;"

Centripetal acceleration a can be expressed: 

"a=\upsilon^2/r;"

So the centripetal force acting on the car :

"F=(m*\upsilon^2)/r;"

Force due to friction:

"f=\mu*N;"

If car is situated in the origin, Force equations at this maximum speed for the car are:

"(m*\upsilon^2)/r=N*sin\theta+\mu*N*cos\theta"

"m*g+\mu*N*sin\theta=N*cos\theta;"

Due to condition, "\mu=0;"

After equations reduce:

"(m*\upsilon^2)/r=N*sin\theta;"

"m*g=N*cos\theta;"

Dive first with second:

"(M*\upsilon^2/r)/mg=(N*sin\theta)/(N*cos\theta);"

As a result we get:

"tan\theta=\upsilon^2/(m*g);"

The final formula is:

"\theta=arctan(\upsilon^2/(m*g));"

"\theta=arctan(35^2/(400*9.8))" "=arctan(0.3125)=17.35;"

Answer: "\theta=17.35."